Circular No. 7905

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002cy IN NGC 1762 L. A. G. Monard, Pretoria, South Africa, reports his discovery of an

apparent supernova on unfiltered CCD images taken with a 0.30-m f/4.2 reflector on May 8.698 (mag 16.7), 12.690 (mag ∼ 16.5), and 13.682 UT (mag 16.1). SN 2002cy is located at α = 5^{h}03^{m}35^{s}.33, δ = +1^{o}34′29′′.5 (equinox 2000.0), which is 30′′ west and 5′′ north of the center of NGC 1762. Nothing is present in this location to limiting mag 19.5 on the Digitized Sky Survey.

SUPERNOVA 2002cz W. M. Wood-Vasey, G. Aldering, and P. Nugent, Lawrence Berkeley

National Laboratory, report the discovery of an apparent supernova (mag 18.8) on NEAT images taken at Palomar (as reported for SN 2002cx on IAUC 7902) on May 9.19 UT, confirmed at mag 18.6 on May 17.3 and 18.8. SN 2002cz is located at α = 13^{h}34^{m}02^{s}.59, δ = −14^{o}49′19′′.7 (equinox 2000.0), which is 9′′ east of the center of the apparent host galaxy. No object is apparent at this position on NEAT images taken on Apr. 30 (limiting mag 20).

SUPERNOVA 2002da Wood-Vasey et al. also report the discovery of an apparent supernova

on NEAT images taken at Palomar on May 16.3 (mag 18.8) and 18.2 UT (mag 18.9). SN 2002da is located at α = 13^{h}24^{m}35^{s}.96, δ = +21^{o}20′08′′.2 (equinox 2000.0), which is 2′′ west of the apparent host galaxy. No object is apparent at this location on images from May 3 (limiting mag 20).

SUPERNOVA 2002cw IN NGC 6700 A. V. Filippenko, University of California, Berkeley, reports that in-

spection of a CCD spectrum (490–820 nm) obtained by M. Brotherton and P. Nandra on May 18 UT with the Mayall 4-m reflector at Kitt Peak reveals that SN 2002cw (cf. IAUC 7902) is of type Ib, a few weeks after the explosion. He I lines with P-Cyg profiles are visible, and there is no obvious Hα emission.

XTE J0929–314

Corrigendum. On IAUC 7900, line 8, for UT read TDB

2002 May 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green